OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 4x – 3y = –6 –x – 5y = –9

OpenStudy (anonymous):

choices are:

OpenStudy (anonymous):

4(–5y + 9) – 3y = –6 4(5y – 9) – 3y = –6 4x – 3(–5y + 9) = –6 4x – 3(5y – 9) = –6

OpenStudy (anonymous):

System of Equations? 4x – 3y = –6 –x – 5y = –9 x=(3y+6)/4 >> x=5y+9 y=(4x+6)/3 y=(-x+9)/5 4(5y+9) - 3y = -6

OpenStudy (anonymous):

Of course, I might have gotten it wrong somewhere... -x = 5y+9 x = -5y+9 4(-5y+9) - 3y = -6

OpenStudy (anonymous):

25x + 10y = 195 substitute eq(1) into eq(2): 25x + 10(15 - x) = 195 25x + 150 - 10x = 195 15x = 45 x = 3 quarters substitute into eq(1): y = 15 - 3 y = 12 dimes Check using eq(2): 25(3) + 10(12) = 195 75 + 120 = 195

OpenStudy (anonymous):

thank you for showing the steps! it's good to know how to do it so im better next time

OpenStudy (anonymous):

wrong question kelly haha